l nulrect, use the hotel guest as the cost object 0l nlrect. For purposes of PB 1-4B. Manufacturing income statement, statement of cost of goodsObj. 4 manufactured y, Several items are omitted from the income statement and cost of goods manufactured statement 00 data for two different companies for the month of December: On Off Company s 195,300 Company Materials inventory, December 1 Materials inventory, December 31 Materials purchased Cost of direct materials used in production Direct labor Factory overhead Total manufacturing costs incurred in December Total manufacturing costs Work in process inventory, December 1 Work in process inventory, December 31 Cost of goods manufactured Finished goods inventory, December 1 Finished goods inventory, December 31 Sales Cost of goods sold Gross profit Operating expenses Net income $ 65,800 ATE 91,140 282,800 317,800 387,800 148,400 577,220 256,060 1,519,000 1,727,320 208,320 973,000 119,000 172,200 1,532,020 269,080 224,000 97,400 ,127,000 ,944,320 1,545,040 117600 164,920 Instructions 1. For both com panies, determine the amounts of the missing items A) through (F), identify ing them by letter. 2. Prepare On Companys statement of cost of goods manufactured for December. 3. Prepare On Companys income statement for December.
Answer #1

1.

On Company

A) Material Consumed = Opening inventory + Purchases-Closing Inventory (A)

317800 = 65800 + 282800 - (A)

Therefore, Material inventory on Dec,31 = $30,800

B) Total Manufacturing Cost (Dec) = Cost of material used + Labour + Factory Overhead

=> 317800 + 387800 + 148400 = $854,000

C) Cost of Goods Manufactured = WIP (DEC-1) + Total manufacturing cost - WIP (Dec-31)

=> 119000 + 973000 - 172700 = 919300

D) Cost of Goods Sold = Finished Goods (DEC-1) + Cost of Goods Manufactured - Finished Goods (Dec-31)

=> 224000 + 919300 - 197400 = 945900

E) Gross Profit = Sales - COGS => 1127000-945900 = $181,100

F) Net Income = Gross Profit - Operating Expenses => 181100-117600 = $63,500

OFF Company

B) Cost of material used = Total manufucaturing cost (Dec) - Labour - Factory OH

=> 1519000-577200-256060 = $685,720

A) Material Purchased = Cost of material used - Material Inventory (Dec 1) + Material Inventory (Dec 31)

=> 685720 - 195300 + 91140 = $581,560

C) WIP (DEC 31) = Total manufacturing cost + WIP (Dec 1) - Cost of Goods Manufactured

=> 1727320 + 208320 - 1532020 = $403,620

D) Finished Goods (Dec 31) = Cost of Goods Manufactured + FG (Dec1) - Cost of Goods Sold

=> 1532020 +269080 - 1545040 = $256,060

E) Gross Profit = Sales - COGS = 1944320 - 1545040 = $399,280

F) Operating Expenses = Gross profit - Net income = 399280 - 164920 = $234,360

2.

Statement of Cost of Goods Manufactured (on Company)
Work in Progress Inventory (Dec 1)       119,000
Add : Total Manufacuring Cost       973,000
Less : Work in Progress Inventory (Dec 31)       172,700
Cost of Goods Manufactured       919,300

3.

Income Statement On Company (December)
Sales    1,127,000
Less : Cost of Goods Sold       945,900
Gross Profit       181,100
Less : Operating Expenses       117,600
Net Income          63,500